OpenStudy (anonymous):

the planck's constant, h, are they two of them? one is 6.262x10^-34 and the other 4.14x10^-15? when do i use them?

OpenStudy (michele_laino):

using the units of measure of the International System of units, or MKS, we have: \[\Large h = 6.62 \times {10^{ - 34}}Joules \times \sec \] If we use the CGS system (Centimeters-Grams-Seconds), we have: \[\Large h = 6.62 \times {10^{ - 27}}erg \times \sec \]

OpenStudy (irishboy123):

your second h is just in eV your first one has a typo, see @Michele_Laino's post